OpenStudy (anonymous):

Derivative of 7/sin(x) + 1/cot(x)

OpenStudy (anonymous):

I don't know what I did wrong but I got the answer as = -7cosx + csc2x

OpenStudy (anonymous):

of am I supposed to use chain rule?

OpenStudy (chrisplusian):

\[\frac{ -7 }{ \cos^2 x } +\sec^2x\] The derivative as the first part is sinx to the negative one power in the numerator times seven. when you bring the negative one down the seven becomes negative and then you have sine x to the negative second power. The derivative of the inside becomes cosx to the negative second which is cos sqared in the denominator. then add that to the derivative of tan because tan is 1 over cot. The derivative of tanx is sec squared x

OpenStudy (chrisplusian):

YOu can probably use other trig identities to simplify this even further but I would have to try it

OpenStudy (chrisplusian):

I made a mistake this is actually\[\frac{ -7 }{ \sin^2 x } * cosx +\sec^2x\] because yes, you should use the chain rule. Sorry I messed upon the first answer

OpenStudy (chrisplusian):

so does that make sense @thuyvy ?

OpenStudy (anonymous):

yes I got it

OpenStudy (chrisplusian):

k :)

OpenStudy (anonymous):

hello, are you still there?

OpenStudy (anonymous):

you can type your formula into wolframalpha

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

They dont show me step by step, can you help

OpenStudy (anonymous):

how did you get tan for the derivative ?

OpenStudy (anonymous):

1/cotx is tanx

OpenStudy (anonymous):

and 1/sinx is cscx

OpenStudy (anonymous):

this will do it, thanks so much :)